Henrich is a single taxpayer. In 2019, his taxable income is 461000. What is his income tax and net investment income tax liability in each of the following alternative scenarios? Use Tax Rate Schedule, Dividends and Capital Gains Tax Rates, Estates and Trusts for reference. c. His 461000 of taxable income includes 61000 of long-term capital gain that is taxed at preferential rates

Income tax-------

Net investment income tax -------

Total tax liability ---------

Please explain how you get the tax on preferencially taxed income don't understand how to get the 15% and the 20% where do those numbers come from?
Answer #1

Part C

Income tax (ordinary income 400000) = 46628.50+(35%*(400000-204100)) = $115193.50 = $115194

Net investment tax = 61000*20% = $12200 (as taxable income is more than $434550)

Total tax liability = $127394

In 2019

For single taxpayers having taxable income $0 - $39,375, preferential tax rate is 0%, $39,376 -$434,550, preferential tax rate is 15%, above $434,550, it is 20%.